Tacoma Hospital has three support departments and four patient services departments. The direct costs to each of the support departments are as follows:

Direct Costs

Cost Driver

General Administration

$2,000,000

Patient Services Revenue

Facilities

$5,000,000

Space (square feet)

Financial Services

$3,000,000

Patient Services Revenue

Total

$10,000,000

The patient services revenue and space utilization (square feet) for each department are as follows:

Patient Services Revenue

Space (square feet)

Routine Care

$30,000,000

400,000

Intensive Care

$4,000,000

40,000

Diagnostic Services

$6,000,000

60,000

Other Services

$10,000,000

100,000

Total

$50,000,000

600,000

Assume that the hospital uses the direct method for cost allocation. Furthermore, the cost driver for general administration and financial services is patient services revenue, while the cost driver for facilities is space utilization.

a. Use an allocation table similar to Exhibit 6.7 to allocate the hospital’s overhead costs to the patient services departments.
Answer #1

a.Allocation of support department overhead costs based on drivers will be as follows:

Routine Care

Intensive Care

Diagnostic Services

Other Services

General Administration

$2,000,000*30/50 = $1,200,000

$2,000,000*4/50 = $160,000

$2,000,000*6/50 = $240,000

$2,000,000*10/50 = $400,000

Facilities

$5,000,000*40/60 = $3,333,333.33

$5,000,000*4/60 = $333,333.33

$5,000,000*6/60 = $500,000

$5,000,000*10/60 = $833,333.33

Financial Services

$3,000,000*30/50 = $1,800,000

$3,000,000*4/50 = $240,000

$3,000,000*6/50 = $360,000

$3,000,000*10/50 = $600,000

Total Overhead Costs

$6,333,333.33

$733,333.33

$1,100,000

$1,833,333.33